# Temperature Optimization for Bayesian Deep Learning

### v1

@@ 4.1 Results
Comparison between SM-PD and TM-PD.
The test LPD across different temperatures for SM-PD and TM-PD are shown in Figure 1; see Appendix K for additional figures. For Concrete, Energy, Naval and MNIST, we observe that TM-PD generally outperforms SM-PD. This is perhaps not surprising, as we expect the former to better account for aleatoric uncertainty. For Concrete and MNIST, the peaks of SM-PD and TM-PD also coincide. Surprisingly, TM-PD does not outperform SM-PD in the CIFAR10 examples.
In terms of the efficacy of our temperature selection method, we find that it can generally recover $\beta$ with a good test LPD. An interesting observation here is that our method tends to select the optimal $\beta$ for TM-PD in the regression examples, and optimal $\beta$ for SM-PD in the classification examples.
We also report MSE (for regression) and accuracy (for classification) to assess the point predictions of the PPDs. As the point predictions are identical across SM-PD and TM-PD by definition, we consolidate the results in Table 1. In general, we observe that SM-PD and TM-PD outperform both SGD and the PPD at $\beta=1$ across both metrics. Furthermore, our method can select temperatures that achieves similar, if not better, performance than that computed from the grid search.
@@
Table 1: MSE (for regression) and accuracy (for classification) of the point predictions of the PPDs at $\beta=1$, $\beta=\hat{\beta}^{*}$ (our method), and the optimal $\beta$ obtained from grid search. We also include the predictions from SGD as a baseline. The results for SM-PD and TM-PD are consolidated, as both produce identical point predictions by definition. The presented values are means $\pm$ standard error across five repetitions, with the best value among the four methods boldfaced. The PPDs at $\beta\neq 1$ generally outperform SGD and the PPD at $\beta=1$, and our method can achieve comparable, if not better, performance than the grid search.
@@
5 CPE through the lens of Bayesian Deep Learning
@@
While the exact interpretation of a ‘well-specified prior’ is debatable, Fortuin et al. (2022) conducted a large-scale experiment to study the effect of tempering under Gaussian (isotropic or with correlated covariance) and heavy-tailed priors (Student’s-t or Laplace) on four metrics: test LPD, accuracy, expected calibration error (Naeini et al., 2015), and out-of-distribution detection accuracy. They observed that $\beta=1$ is indeed optimal for test LPD and accuracy when using a heavy-tailed prior. However, the standard posteriors derived from heavy-tailed priors also tend to underperform compared to tempered posteriors derived from Gaussian priors. Notably, according to their experiments, warmer temperatures tend to perform better for expected calibration error, and there is no general trend in optimal $\beta$ for out-of-distribution detection accuracy. Therefore, their work suggests that the optimal $\beta$ is not only dependent on the truth-model-prior triplet but also on the evaluation metric.
@@
6 CPE through the lens of Generalized Bayes
In GB, it is a standard practice to tune $\beta$ to improve the utility of the tempered posterior (Zhang, 2006a; Grünwald, 2012; Bissiri et al., 2016). Specifically, the likelihood is frequently (but not always) tempered to a warmer temperature, which contrasts with the cold temperature used in BDL. In this section, we review some recent developments in GB to clarify the role of tempering in GB and this apparent contradiction.
@@ 6 CPE through the lens of Generalized Bayes
It is easy to verify that the inequality holds at $\beta=1$ for well-specified models, where $q(x,y)=p(y|x,\theta_{\dagger})q(x)$. The same cannot be said for misspecified models in general. However, Grünwald (2012) shows that this inequality can still hold for many misspecified models for some warm temperatures $\beta\leq\beta_{\text{critical}}<1$, and proposed the SafeBayes algorithm to determine $\beta_{\text{critical}}$. Therefore, SafeBayes is presented as a tool to achieve posterior concentration on $\theta_{\dagger}$.
However, posterior concentration is not always desirable in the context of maximizing test LPD under model misspecification. Note that taking $\beta\leq\beta_{\text{critical}}$ results in the corresponding PPD concentrating on the ‘plug-in’ predictive density $p(y|x,\theta_{\dagger})$ as $n\to\infty$. This may be undesirable if one prioritizes the test LPD of the PPD, since the plug-in $p(y|x,\theta_{\dagger})$ may not exhibit good test LPD. For example, when the misspecified model is non-convex, the PPD can lie outside the model family and be closer to the truth than the plug-in $p(y|x,\theta_{{\dagger}})$. Therefore, avoiding posterior concentration in this situation can actually result in a PPD with a higher test LPD.
@@
Concurrent to our work, McLatchie et al. (2024) analyzed the role of $\beta$ [^4] In McLatchie et al. (2024), the learning rate and, confusingly, temperature refer to $\beta$ in this work. in improving test LPD theoretically. They concluded that, in the moderately large $n$ regime and assuming posterior concentration at $\theta_{{\dagger}}$, the test LPD shows little improvement once $\beta$ becomes sufficiently large. Although their motivation aligns with ours, their theoretical results do not extend to singular models (e.g., neural networks) or over-parameterized models (e.g., ResNet20 on CIFAR10).
6.1 Limitations of the GB framework for analyzing BDL models
@@ 8 Reproducibility Statement
The experiments details are summarized in the introduction of Section 4. Proof of Lemma 3.1 is given in Appendix E. SGMCMC is described in Appendix H. Diagnostic check of the SGMCMC is given in Appendix I. Implementation details of our proposed method is given in Appendix C. The set of hyperparameters to reproduce the experiments is given in Appendix J. Computational environment and packages are given in Section J.1. Evaluation metrics are defined in Appendix F. Model, prior and datasets are described in Appendix G. Source code is published on https://anonymous.4open.science/r/tempered-posteriors-106E.
@@ 9 Acknowledgment
We would like to thank David Frazier and Dino Sejdinovic for helpful discussions. This research is supported by Cloud TPUs from Google’s TPU Research Cloud. KN is supported by the Australian Government’s Research Training Program and a joint top-up scholarship from the Statistical Society of Australia and the Australian Bureau of Statistics. LH is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144).
@@ Appendix A Optimal temperature of a Gaussian Toy Model
$=\mathcal{N}\left(\mu\middle|\bar{x}\left(\frac{\sigma^{2}T}{n\sigma^{2}_{p}}+1\right)^{-1},\sigma^{2}_{post}\right),$
@@ Appendix D Variants of the Temperature Selection Method
In this section, we compare a variant of (7) by maximizing
$\hat{\theta}^{*},\hat{\beta}^{*}\coloneq\operatorname*{arg\,max}_{\theta,\beta}\frac{1}{n}\sum_{(x,y)\in\mathcal{D}}[\log p(y|x,\theta,\beta)+\log p(\theta)].$ (9)
We shall call this the maximum-a-posteriori (MAP) method, as $\theta$ is now being constrained by a prior $p(\theta)$. This is in contrast to the maximum likelihood (MLE) method as introduced in (7). We compare both MAP and MLE on the CIFAR10 experiments and show the achieved test LPD and accuracy in Table 2. In general, we find that MAP and MLE perform similarly well.
@@
Table 2: Comparison between strategies to select $\beta$ on the CIFAR10 experiments. Reported values are mean $\pm$ standard error across five repetitions. All values are evaluated on a test set. The accuracy and LPD are all within one standard error of difference between MAP and MLE.
Data Strategy Accuracy $\uparrow$ LPD $\uparrow$ $\hat{\beta}^{*}$
CIFAR10 MAP $89.89\pm 0.25$ $-0.321\pm 0.010$ $10.54\pm 9.60$
@@ I.2 Rank-normalized split-$\hat{R}$
In this section, we provide a detailed explanation of the hyperparameters for SGD and SGMCMC. The values are summarized in Table 6.
@@
Appendix K Extra Figures From the Main Experiment
In this section, we provides extra figures to complement the results shown in the main text.
We first shows the validation LPD plotted against the temperature in Figure 3, and find that our method can recover temperatures in the high performing regions. We further provide a better visualization of Figure 1 by separating the test LPD computed with SM-PD (3) and TM-PD (8) into Figure 4 and Figure 5 respectively. A SGD reference is also included.
Lastly, we shows the performance of the point predictions across different temperatures in Figure 6, to complement Table 1. Note that the point predictions are the same across SM-PD and TM-PD. We observe that the predictions from tempered posteriors generally outperforms SGD.
@@
The classical Bayesian rule (1) updating prior belief $p(\theta)$ to the posterior belief $p(\theta|\mathcal{D})$ arises from Bayes’ Theorem. An implicit assumption in this procedure for optimal performance is that the model is correctly specified: there exists some $\theta_{0}$ such that $p(y|x,\theta_{0})=q(y|x)$ for all $(x,y)$. The adage that “all models are wrong” Box (1976) highlights that this assumption should not be expected to hold. Consequently, it does not make sense to infer on a parameter $\theta$ that lacks a connection to the true data-generation distribution.

### v2

@@ 4.1 Results
Optimal temperature of SM-PD and TM-PD across models and datasets.
The test LPD across different temperatures for SM-PD and TM-PD are shown in Figure 1; see Appendix K for additional figures and Table 7 and Table 8 for tabular versions of the results. For Concrete, Energy, Naval and MNIST, we observe that TM-PD generally outperforms SM-PD. This is perhaps not surprising, as we expect the former to better account for aleatoric uncertainty. For Concrete and MNIST, the peaks of SM-PD and TM-PD also coincide. Surprisingly, TM-PD does not outperform SM-PD in the CIFAR10 examples.
In terms of the efficacy of our temperature selection method, we find that it can generally recover $\beta$ with a good test LPD. An interesting observation here is that our method tends to select the optimal $\beta$ for TM-PD in the regression examples, and optimal $\beta$ for SM-PD in the classification examples. Moreover, as shown in Figure 1, the optimal $\beta$ for TM-PD and SM-PD for a given model are usually different. Therefore, we expect $\hat{\beta}^{*}$, which is computed irrespective of the construction of the PPD, may only work for either TM-PD or SM-PD, but not both at the same time.
We also report MSE (for regression) and accuracy (for classification) to assess the point predictions of the PPDs. As the point predictions are identical across SM-PD and TM-PD by definition, we consolidate the results in Table 1. In general, we observe that SM-PD and TM-PD outperform both SGD and the PPD at $\beta=1$ across both metrics. Furthermore, our method can select temperatures that achieves performance comparable to that of the grid search.
@@
Table 1: MSE (for regression) and accuracy (for classification) of the point predictions of the PPDs at $\beta=1$, $\beta=\hat{\beta}^{*}$ (our method), and the optimal $\beta$ obtained from grid search. We also include the predictions from SGD as a baseline. The results for SM-PD and TM-PD are consolidated, as both produce identical point predictions by definition. The presented values are means $\pm$ standard error across five repetitions, with the best value among the four methods boldfaced.
@@
Computation time
Our approach requires a single SGD run to compute the optimal $\beta$ and a SGMCMC run to generate the PPD, whereas the grid search requires 9 SGMCMC runs — one for each temperature in the grid. The wall-clock times for the main experiment are reported in Table 9 of Appendix K. Overall, our method is 4 times faster than the grid search for the smaller regression models and 8 times faster for the larger classification models.
@@
5 Related Work: CPE through the Lens of Bayesian Deep Learning
@@
While the exact interpretation of a ‘well-specified prior’ is debatable, Fortuin et al. (2022) conducted a large-scale experiment to study the effect of tempering under Gaussian (isotropic or with correlated covariance) and heavy-tailed priors (Student’s-t or Laplace) on four metrics: test LPD, accuracy, expected calibration error (Naeini et al., 2015), and out-of-distribution detection accuracy. They observed that $\beta=1$ is indeed optimal for test LPD and accuracy when using a heavy-tailed prior. However, the standard posteriors derived from heavy-tailed priors also tend to underperform compared to tempered posteriors derived from Gaussian priors. Notably, warmer temperatures tend to lower expected calibration error, and there is no general trend in optimal $\beta$ for out-of-distribution detection accuracy. Therefore, their work suggests that the optimal $\beta$ is not only dependent on the truth-model-prior triplet but also on the evaluation metric.
@@
6 Related Work: CPE through the Lens of Generalized Bayes
In GB, it is common to tune $\beta$ to improve the utility of the tempered posterior (Zhang, 2006a; Grünwald, 2012; Bissiri et al., 2016). Specifically, the likelihood is frequently (but not always) tempered to a warmer temperature, which contrasts with the cold temperature used in BDL. In this section, we review some recent developments in GB to clarify the role of tempering in GB and this apparent contradiction.
@@ 6 Related Work: CPE through the Lens of Generalized Bayes
This inequality holds at $\beta=1$ for well-specified models, where $q(x,y)=p(y|x,\theta_{\dagger})q(x)$, but the same cannot be said for misspecified models in general. Grünwald (2012) shows that this inequality can still hold for many misspecified models for some warm temperatures $\beta\leq\beta_{\text{critical}}<1$, and proposed the SafeBayes algorithm to determine $\beta_{\text{critical}}$. Therefore, SafeBayes is presented as a tool to achieve posterior concentration on $\theta_{\dagger}$.
However, posterior concentration is not always desirable in the context of maximizing test LPD under model misspecification. By taking $\beta\leq\beta_{\text{critical}}$, the corresponding PPD will concentrate on the ‘plug-in’ predictive density $p(y|x,\theta_{\dagger})$ as $n\to\infty$. This may be undesirable, since $p(y|x,\theta_{\dagger})$ may not exhibit good test LPD. For example, when the misspecified model is non-convex, the PPD can lie outside the model family and be closer to the truth than $p(y|x,\theta_{{\dagger}})$. Therefore, avoiding posterior concentration in this situation can actually result in a PPD with a higher test LPD.
@@
Concurrent to our work, McLatchie et al. (2024) analyzed the role of $\beta$ [^4] In McLatchie et al. (2024), the learning rate and, confusingly, temperature refer to $\beta$ in this work. in improving test LPD theoretically. They concluded that, in the moderately large $n$ regime and assuming posterior concentration at $\theta_{{\dagger}}$, the test LPD shows little improvement once $\beta$ becomes sufficiently large. Although their motivation aligns with ours, their theoretical results do not extend to neural networks or over-parameterized models (e.g., ResNet20 on CIFAR10).
6.1 Limitations of Generalized Bayes for Analyzing BDL Models
@@ 8 Reproducibility Statement
The experiments details are summarized in the introduction of Section 4. Proof of Lemma 3.1 is given in Appendix E. SGMCMC is described in Appendix H. Diagnostic check of the SGMCMC is given in Appendix I. Implementation details of our proposed method is given in Appendix C. The set of hyperparameters to reproduce the experiments is given in Appendix J. Computational environment and packages are given in Section J.1. Evaluation metrics are defined in Appendix F. Model, prior and datasets are described in Appendix G. Source code is published on https://github.com/weiyaw/tempered-posteriors.
@@ 9 Acknowledgment
We would like to thank David Frazier, Dino Sejdinovic and the anonymous reviewers for their helpful discussions. This research was supported by Cloud TPUs from Google’s TPU Research Cloud. K.N. is supported by the Australian Government’s Research Training Program and a joint top-up scholarship from the Statistical Society of Australia and the Australian Bureau of Statistics. L.H. is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144).
@@ Appendix A Optimal Temperature of a Gaussian Toy Model
$=\mathcal{N}\left(\mu\middle|\bar{x}\left(\frac{\sigma^{2}}{n\sigma^{2}_{p}}+1\right)^{-1},\sigma^{2}_{post}\right),$
@@ Appendix D Variants of the Temperature Selection Method
In this section, we compare a variant of (7) by solving
$\operatorname*{arg\,max}_{\theta,\beta}\frac{1}{n}\sum_{(x,y)\in\mathcal{D}}[\log p(y|x,\theta,\beta)+\log p(\theta)].$ (9)
We refer to this approach as the maximum-a-posteriori (MAP) method, as $\theta$ is now constrained by a prior $p(\theta)$. This contrasts with the maximum likelihood (MLE) method introduced in (7). Additionally, we compare our method to the popular temperature scaling approach proposed in Guo et al. (2017). In their method, $\theta$ in the objective (7) is fixed to a solution obtained from the standard training workflow (e.g., $\theta^{*}_{\text{SGD}}$), and the temperature $\beta$ is optimized by solving
$\operatorname*{arg\,max}_{\beta}\frac{1}{n}\sum_{(x,y)\in\mathcal{D_{\mathrm{valid}}}}\log p(y|x,\theta^{*}_{\text{SGD}},\beta)$
on a validation set. In contrast, both MLE and MAP jointly optimizes $\theta$ and $\beta$ within a single SGD run.
We compare the methods on the CIFAR-10 experiments, reporting the test LPD and accuracy in Table 2. The PPDs are constructed from SM-PD (3). In general, MAP and MLE yield similar $\beta$ values, while the $\beta$ obtained from Guo et al. (2017) differs substantially. Moreover, Guo et al. (2017) generally underperforms relative to MAP and MLE when the CPE is most pronounced, such as when data augmentation is enabled. This outcome is perhaps unsurprising, as Guo et al. (2017) is designed for computing a well-calibrated tempered model $p(y|x,\theta,\beta)$, whereas our method targets the PPD $p_{\beta}(y|x,\mathcal{D})$.
The results also suggest that optimizing $\theta$ and $\beta$ together is critical for determining a good $\beta$. We conjecture that $p(y|x,\theta^{*},\beta^{*})$, with both $\theta^{*}$ and $\beta^{*}$ obtained through our procedure, provides a reasonable approximation of the PPD with the highest test LPD. In contrast, fixing $\theta$ to a predetermined value (e.g., $\theta^{*}_{\text{SGD}}$) limits the search space and leads to a poor approximation of the PPD.
@@
Table 2: Comparison between methods to select $\beta$ on the CIFAR10 experiments. Reported values are mean $\pm$ standard error across five repetitions. All values are evaluated on a test set. The accuracy and LPD are all within one standard error of difference between MAP and MLE.
Data Method Accuracy $\uparrow$ LPD $\uparrow$ $\hat{\beta}^{*}$
CIFAR10 Guo et al. (2017) $89.86\pm 0.2$ $-0.317\pm 0.003$ $1.75\pm 0.18$
MAP $89.89\pm 0.25$ $-0.321\pm 0.010$ $10.54\pm 9.60$
@@ I.2 Rank-normalized Split-$\hat{R}$
In this section, we provide a detailed explanation of the hyperparameters for SGD and SGMCMC. The values are summarized in Table 6. Our source code is available on https://github.com/weiyaw/tempered-posteriors.
@@
Appendix K Extra Figures and Tables From the Main Experiment
In this section, we provide additional figures and tables to complement the results presented in the main text.
We first show the validation LPD plotted against the temperature in Figure 3 and find that our method can recover temperatures in the high-performing regions. We further improve the visualization of Figure 1 by separating the test LPD computed with SM-PD (3) and TM-PD (8) into Figure 4 and Figure 5, respectively. A tabular version of the results is also provided in Table 7 and Table 8, with an SGD reference included in all the results.
We also show the performance of point predictions across different temperatures in Figure 6, complementing Table 1. Note that the point predictions are identical for SM-PD and TM-PD. We observe that predictions from tempered posteriors generally outperform SGD.
Finally, we show the wall-clock time comparison across the methods in Table 9. Overall, our method is 4 times faster than the grid search for the smaller regression models and 8 times faster for the larger classification models.
Table 7: Test LPD of the regression models. The values presented are means $\pm$ standard errors across five repetitions, with the best value among the four methods highlighted in bold. TM-PDs generally have better performance for the regression models. Higher LPD indicates better performance.
Method Concrete Energy Naval
SM-PD SGD $-3.904\pm 1.000$ $1.299\pm 0.011$ $1.376\pm 0.003$
$\beta=1$ $-0.506\pm 0.255$ $1.221\pm 0.005$ $1.376\pm 0.000$
$\beta=\hat{\beta}^{*}$ $\mathbf{-0.186\pm 0.049}$ $1.287\pm 0.009$ $\mathbf{1.382\pm 0.000}$
Grid $-0.216\pm 0.121$ $\mathbf{1.312\pm 0.003}$ $\mathbf{1.382\pm 0.001}$
TM-PD SGD $-0.319\pm 0.076$ $1.75\pm 0.054$ $2.95\pm 0.162$
$\beta=1$ $-0.506\pm 0.255$ $1.22\pm 0.005$ $1.38\pm 0.000$
$\beta=\hat{\beta}^{*}$ $-0.13\pm 0.069$ $1.78\pm 0.063$ $3.16\pm 0.106$
Grid $\mathbf{0.019\pm 0.035}$ $\mathbf{1.85\pm 0.017}$ $\mathbf{3.71\pm 0.194}$
Table 8: Test LPD of the classification models. The presented values are the means $\pm$ standard error across five repetitions, with the best value among the four methods boldfaced. SM-PDs tend to perform better on CIFAR-10, while SM-PD and TM-PD have similar performance on MNIST. Higher LPD indicates better performance.
Method MNIST CIFAR10 CIFAR10 (DA)
SM-PD SGD $-0.076\pm 0.032$ $-1.138\pm 0.306$ $-1.234\pm 0.176$
$\beta=1$ $-0.021\pm 0.000$ $-0.333\pm 0.001$ $-0.343\pm 0.002$
$\beta=\hat{\beta}^{*}$ $\mathbf{-0.018\pm 0.000}$ $\mathbf{-0.317\pm 0.006}$ $-0.232\pm 0.007$
Grid $\mathbf{-0.018\pm 0.000}$ $-0.319\pm 0.003$ $\mathbf{-0.229\pm 0.004}$
TM-PD SGD $-0.030\pm 0.003$ $-0.55\pm 0.036$ $-0.284\pm 0.004$
$\beta=1$ $-0.021\pm 0.000$ $\mathbf{-0.333\pm 0.001}$ $-0.343\pm 0.002$
$\beta=\hat{\beta}^{*}$ $\mathbf{-0.017\pm 0.000}$ $-0.719\pm 0.443$ $-0.811\pm 0.221$
Grid $\mathbf{-0.017\pm 0.001}$ $\mathbf{-0.333\pm 0.001}$ $\mathbf{-0.279\pm 0.007}$
@@
Table 9: Wall-clock time comparison across methods. Times are averaged over five repetitions.
Method Wall-clock time (minutes)
Concrete Energy Naval MNIST CIFAR10 CIFAR10 (DA)
SGD 0.17 0.16 0.78 1.93 27.90 34.84
$\beta=1$ 0.20 0.19 1.40 18.83 260.52 327.12
$\beta=\hat{\beta}^{*}$ (Ours) 0.37 0.36 2.19 20.77 288.42 361.96
Grid 1.76 1.74 12.64 169.51 2344.15 2944.01
@@
The classical Bayesian rule (1) updating prior belief $p(\theta)$ to the posterior belief $p(\theta|\mathcal{D})$ arises from Bayes’ Theorem. An implicit assumption in this procedure for optimal performance is that the model is correctly specified: there exists some $\theta_{0}$ such that $p(y|x,\theta_{0})=q(y|x)$ for all $(x,y)$. The adage that “all models are wrong” (Box, 1976) highlights that this assumption should not be expected to hold. Consequently, it does not make sense to infer on a parameter $\theta$ that lacks a connection to the true data-generation distribution.